OpenStudy (anonymous):

Swine flu is attacking porkopolis. the function S(t) = 9t - 4 determines how many people have swine where t = time in days and S = number of people in thousands.

OpenStudy (dtan5457):

What's the question though?

OpenStudy (dtan5457):

Reread what you posted, you aren't asking anything

OpenStudy (anonymous):

I edited the problem and added the given equation. I have to find S(4), find t when S(t)=23,

OpenStudy (sleepyhead314):

S(4) means to plug in 4 for t and simplify so 9(4) - 4 = ? then S(t) = 23 becomes 23 = 9t - 4 and solve for that t from there

OpenStudy (dtan5457):

S(t) = 9t - 4 s(4)= 9(4)-4=s(t)(4) 2nd one s=9(23)-4 remember to put your answer by 1000

OpenStudy (anonymous):

thank you sleepyhead314, I should be able to to work it from here dtan5457 thanks, i will remember.

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :3